Air handling unit HVAC Analysis

known ambient air characteristics:
Relative humidity = 55%
air is 79% n2 and 21% o2
I have the mole fractions and the mass fractions entering the fan
the fan has a inlet pressure = atmospheric = 0 psig = 14.7 psia

fan characteristics:
fan duct = 4ft diameter
air mixture velocity = 14ft/second

the fan outlet pressure is 0.9 psig = 15.6 psia

with a known fan fluid efficiency of 60%
and know CV and CP values

how can I calculate the increase in Temperature of the air mixture using ideal gas laws? Thanks!
Answer #1

Inlet pressure P = 147 paia & 101.312 R Rareal fan duct = 4 ft diameter 8= 2 ft ration ( air mixture velocity = 14 ft /sec Ou0.6 n = 60% = put in equ Job I-a 1.061) s y = 1.4 loc CP - 1.4 I Cp = lover Now Pv=nRT volume = IB-6283 P - 14.7 psia m=55 Ri hope it will please comment for further questions

Please rate positively